### v5

# SCMM: Calibrating Cross-modal Representations for Text-Based Person Search

Abstract
Text-Based Person Search (TBPS) is a crucial task in the Internet of Things (IoT) domain that enables accurate retrieval of target individuals from large-scale galleries with only given textual caption. For cross-modal TBPS tasks, it is critical to obtain well-distributed representation in the common embedding space to reduce the inter-modal gap. Furthermore, learning detailed image-text correspondences is essential to discriminate similar targets and enable fine-grained search. To address these challenges, we present a simple yet effective method named Sew Calibration and Masked Modeling (SCMM) that calibrates cross-modal representations by learning compact and well-aligned embeddings. SCMM introduces two novel losses for fine-grained cross-modal representations: Sew calibration loss that aligns image and text features based on textual caption quality, and Masked Caption Modeling (MCM) loss that establishes detailed relationships between textual and visual parts. This dual-pronged strategy enhances feature alignment and cross-modal correspondences, enabling accurate distinction of similar individuals while maintaining a streamlined dual-encoder architecture for real-time inference, which is essential for resource-limited sensors and IoT systems. Extensive experiments on three popular TBPS benchmarks demonstrate the superiority of SCMM, achieving 73.81%, 64.25%, and 57.35% Rank-1 accuracy on CUHK-PEDES, ICFG-PEDES, and RSTPReID, respectively.
Index Terms:
Text-Based Person Search, Internet of Things (IoT), Person Re-identification, Cross-modal Learning, Metric Learning
I Introduction
With the rapid growth of Internet of Things (IoT) devices and the increasing need for intelligent analysis of visual data, person re-identification (Re-ID) has emerged as a critical task in IoT and multimedia applications [1, 2, 3]. Re-ID aims to identify a target person across non-overlapping camera views, enabling applications such as tracking individuals in crowded scenes and enhancing security measures [4, 5]. However, traditional Re-ID relies on the availability of query images, which may not always be feasible in real-world IoT scenarios. For example, a security guard responding to a witness report describing a suspect’s appearance; in such cases, relying solely on textual descriptions for person retrieval becomes crucial [6, 7]. To address this new scenario, Text-Based Person Search (TBPS) [8] has recently gained increasing interest, which aims to retrieve images of individuals from a large-scale gallery via textual captions as query [9, 10]. In IoT environments, TBPS can be integrated with various sensors and edge server units to enhance real-time processing capabilities and reduce latency. For instance, surveillance cameras  [10] equipped with edge computing units can preprocess video streams and extract textual descriptions of individuals, which are then used for efficient person search across the network. This integration is essential for applications in smart cities and industrial settings  [11], where timely and accurate person identification is crucial for maintaining security and operational efficiency.
Figure 1: Illustration of the motivation behind our method. In cross-modal tasks, a compact and well-aligned image-text feature distribution in the shared embedding space is crucial for bridging the inter-modal gap. Additionally, capturing fine-grained image-text correspondences is equally vital to distinguish between similar individuals for text-based person searching.
Compared to Re-ID, the main task of TBPS is to learn fine-grained cross-modal representations between visual and textual modalities. As shown in Fig. 1, cross-modal representations have two key characteristics that contribute to searching for a person from images: closely aligned cross-modal representations and fine-grained information correspondence. First, closely aligned cross-modal representations can reduce the inter-modal gap, making it easier to locate specific persons with textual captions. Second, detailed correspondence is essential for TBPS. In some cases, fine-grained cross-modal information is necessary to discriminate between two similar persons.
In the TBPS task, numerous approaches have been proposed from two perspectives. Firstly, some methods [12, 13, 14] employ only two encoders and align the two modal representations using a symmetric loss. Although simple, the feature alignment is limited in their methods. The other methods [15, 16] utilize multi-modal models with transformer-based [17] cross-attention to improve cross-modal feature alignment and interaction. However, such methods often require high computational costs due to the fusion of all possible image-text pairs during inference, hindering their efficiency in real-time applications. Secondly, to obtain robust fine-grained features, recent TBPS works have designed multi-level [18, 19, 20, 21], multi-granularity [22, 23] matching strategies, and specific attention modules [10, 24, 25]. These methods rely on the image-text backbone to provide fine-grained features. While these methods provide fine-grained features, they have complex model architectures and costly computations. Furthermore, the fine-grained features they produce are limited and hinder performance boost.
To address these issues, we propose a simple yet effective method called SCMM for calibrating cross-modal representations in TBPS. SCMM consists of only a dual-encoder, making it simple and cost-effective without requiring complex interaction modules or extra multi-level branches, which allows for high-speed inference. In addition, we propose two novel training losses to calibrate cross-modal representations. The first is a Sew calibration loss, which takes the quality of the text description as guidance and aligns features between the textual and visual modalities. It also pushes negative sample pairs apart and pulls positive sample pairs together across the two modalities. Next, we propose a Masking Caption Modeling (MCM) loss to obtain more fine-grained and generic correspondence. This loss uses a masked caption prediction task to establish detailed relationships between text parts and image parts. The operation is implemented through a cross-modal decoder that is discarded at the inference stage, avoiding extra computation costs. To demonstrate the effectiveness of SCMM, we evaluated its performance on three popular benchmarks: CUHK-PEDES [10], ICFG-PEDES [19], and RSTPReID [26]. Our model surpasses the previous state-of-the-art (SOTA) methods and demonstrates impressive performance. Moreover, we conducted extensive experiments to validate each component of our method. Overall, our major contributions can be summarised as follows:
  • We introduce an effective and scalable framework to learn and calibrate cross-modal representations for text-based person search. Our framework utilizes a dual-encoder and an auxiliary cross-modal decoder to achieve efficient and high-speed inference.
  • We propose two novel losses, in which the Sew calibration loss aligns fine-grained features between image and text modalities, as well as the MCM loss establishes detailed relationships between vision modality and textual modality.
  • Extensive experiments demonstrate the superiority of our framework. Our method achieves new state-of-the-art on three popular benchmarks: CUHK-PEDES, ICFG-PEDES, and RSTPReID, which reaches 73.81%, 64.25%, and 57.35% on the Rank1, respectively.
Figure 2: Overview of our proposed SCMM. The framework consists of a dual-encoder for extracting image-text features and calibrating cross-modal representations with the Sew Calibration loss. We also include a decoder for performing cross-modal interaction with the task-driven Mask Caption Modeling. At the inference stage, we only utilize the classification (CLS) tokens from the dual-encoder to implement similarity search.
The structure of the remainder of this paper is outlined as follows. Sec. II reviews the related work on text-based person search, metric learning, and masked language modeling. Sec. III introduces the proposed SCMM method, including the Sew Calibration Loss, Masking Caption Modeling Loss, and the overall framework. Sec. IV presents the experimental setup, datasets, implementation details, and comprehensive experimental results, including comparison with state-of-the-art methods, ablation studies on different components, and qualitative analysis with visualization of attention maps and feature distributions. Sec. V concludes the paper with a summary of contributions and future research directions.
II Related Work
II-A Text-Based Person Search
Text-based person search Textis was first introduced by [10], which identifies person images in a gallery only using a textual query. Early works utilized pre-task methods to obtain external cues such as person segmentation [18] and human body landmarks [27]. In recent years, end-to-end frameworks based on attention mechanisms [10, 28, 24, 21, 23, 25] have become prevailing. Cross-modal attention is critical for performing image-text interaction. Existing methods can be broadly classified into attention-explicit and attention-implicit. Specifically, attention-explicit methods [10, 24, 25] design specific attention modules according to multi-granularity and multi-level strategies. For example, NAFS [28] conducts cross-modal alignments over full-scale features with a contextual non-local attention module. CFine [24] utilizes cross-attention for multi-grained global feature learning, it achieves impressive results on three benchmarks with knowledge transfer from the CLIP [29] model. In contrast attention-implicit methods [21, 23] utilize transformer-based models with shared parameters to align cross-modal semantics implicitly. SafaNet [21] introduces a semantic-aligned feature aggregation network. It utilizes a cross-modal parameter-sharing multi-head attention module following the backbone to enhance the extracted image-text representations. However, compared to performing in-depth cross attention with a task-driven approach, the existing methods do not implement enough cross-interaction with generalized performance.
II-B Metric Learning
Initially, metric learning used $L_{2}$ distance as the metric, with the goal is to minimize the $L_{2}$ distance between samples of the same class. Some $L_{2}$-based metric learning methods include Siamese Networks [30] and Triplet Networks [31]. With the development of deep learning, researchers started using softmax-based loss functions in metric learning to learn a more discriminative distance metric. Additionally, increasing the margin between classes is an intuitive approach to learn a better metric space. L-Softmax [32] introduced the concept of margin on softmax function for the first time. The widely used CosFace [33] proposed large-margin cosine loss to learn highly discriminative deep features for face recognition. Circle loss [34] proposed a simple loss based on a unified loss for metric learning and classification. Recently, some works introduced adaptive margin into marginal loss [35, 36, 11]. They usually learn image quality implicitly and adjust the margin accordingly. In single-modal representation learning, they usually give large margins to high-quality samples for hard mining. Compared to them, we try to solve a cross-modal matching problem where samples from two modalities have different information volume. We give greater tolerance to less informative samples in TBPS.
II-C Masked Language Modeling
MLM is a highly effective method for pre-training language models [37] by randomly selecting a certain percentage of words from the input sentence and then predicting the masked words based on the context of other words. Many cross-modal pre-training models have utilized MLM in their methods [38, 15]. For example, the work in [15] combines MLM with contrastive loss in the framework, which achieved impressive performance in their cross-modal tasks. The success of MLM in BERT [37] has proven its ability to adapt well to various downstream tasks, leading to generalized performance. In our fine-grained framework, the task-driven decoder utilizing masked caption modeling can facilitate generic cross-modal learning.
III Methods
Formally, given a set of images with corresponding captions, denoted as $X=\{(I_{i},T_{i})\}^{N}_{i=1}$. Each image $I_{i}$ and its description text $T_{i}$ is associated with a person ID $y_{i}$. Text-based person search aims to retrieve the most relevant Rank $k$ (e.g., $k=1,5,10$) person images efficiently from a large-scale gallery with a textual caption. To solve this task, we propose a simple yet effective method, as shown in Fig. 2.
We use ViT [39] and BERT [37] as the image-text encoders in our dual-encoder backbone. The image encoder takes image patches from $I_{i}$ along with a vision classification (CLS) token as input. It outputs an image feature sequence ${v_{i}}$ and a vision CLS token embedding $v_{i}^{c}$. Similarly, the text encoder obtains a text feature sequence ${t_{i}}$ and a text CLS token embedding $t_{i}^{c}$ for caption $T_{i}$, following previous works [28, 24].
III-A Sew Calibration Loss with Constraints
In the TBPS task, closely aligning cross-modal representations is crucial for effectively finding specific persons with textual captions. To address the heterogeneity between modalities, we propose a Sew calibration loss that pushes each modality to a common space.
As shown in Fig. 3(a), in single modality settings, a triplet distance constraint is widely utilized when embeddings are from a single modality distribution. In intra-class samples, the embeddings are pulled together, while inter-class samples are pushed away. In the cross-modal setting, we expect to impose constraints from both sides of the two embedding distributions. For example, in one-direction retrieval, we first need to align the features of a ”perfect pair,” i.e., $(I_{i},T_{i})$. The shortest distance from the image embedding ${v_{i}}$ to the text embedding should be ${t_{i}}$. This is because it is a perfect matching pair in person re-identification, and no other text feature will have a shorter distance (Eq. 1). We set the perfect pair as an image-text anchor $(A_{img},A_{txt})$. Next, we impose another constraint between the image anchor $A_{img}$ and its corresponding positive text samples $P_{txt}(\mathds{1}(y_{i}=y_{j},i\neq j))$ and negative text samples $N_{txt}$ (Eq. 2). For the other direction, we put symmetric constraints on $A_{txt}$ and $P_{img}$, $N_{img}$. Fig. 3(b) shows the proposed constraints for cross modality. Forces from both sides act like a seam to pull the two distributions together.
Figure 3: Illustration of Sew calibration loss. The constraints are different between single-modal and cross-modal matching. $(A_{img},A_{txt})$ denotes anchors in image-text feature distribution, while $(P_{img},P_{txt})$ and $(N_{img},N_{txt})$ denote positive and negative sample pairs, respectively. The Sew calibration loss pushes negative sample pairs and pulls positive sample pairs, stitching cross-modal key information like a seam.
Take the image side as an example, the $L_{2}$ distance constraints are shown as follows:
$D(A_{img},A_{txt})+\mathcal{M}_{1}<D(A_{img},P_{txt}),$ (1)
$D(A_{img},P_{txt})+\mathcal{M}_{2}<D(A_{img},N_{txt}),$ (2)
where $D$ denotes $L_{2}$ distance. $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are two margins, $\mathcal{M}_{1}<\mathcal{M}_{2}$. With a bi-directional margin, each modal feature of the same pedestrian target is compressed compactly, making decision boundaries clearer. We then relax the constraints to:
$0.5D^{2}(A_{img},A_{txt})+\mathcal{M}_{1}<0.5D^{2}(A_{img},P_{txt}),$ (3)
$0.5D^{2}(A_{img},P_{txt})+\mathcal{M}_{2}<0.5D^{2}(A_{img},N_{txt}).$ (4)
Subsequently, we change the above pairwise constraints into soft forms for better convergence following [34].
$\mathcal{L}^{Pull}_{match}=log\big{[}1+\sum\limits^{K}_{k=1}exp(\alpha(\bar{v}_{i}^{c}\bar{t}_{k}^{c}-\bar{v}_{i}^{c}\bar{t}_{i}^{c}+\mathcal{M}_{1}))\big{]},$ (5)
$\mathcal{L}^{Push}_{match}=log\big{[}1+\sum\limits^{K}_{k=1}\sum\limits^{J}_{j=1}exp(\alpha(\bar{v}_{i}^{c}\bar{t}_{j}^{c}-\bar{v}_{i}^{c}\bar{t}_{k}^{c}+\mathcal{M}_{2}))\big{]},$ (6)
where $\bar{v}^{c}$ and $\bar{t}^{c}$ are CLS token features after normalizing, $K$ and $J$ denote the number of positive and negative samples in this batch, respectively. $\alpha$ is a scale parameter. As a result, the image-to-text matching part of Sew calibration loss is formulated as below:
$\mathcal{L}^{I2T}_{match}=\mathcal{L}^{Pull}_{match}+\mathcal{L}^{Push}_{match}.$ (7)
The constraints in Eq. 2 can also be used to impose a classification loss in a similar way. As there is no difference for perfect positive samples in the classification task, we omit to constrain Eq. 1. Formally, the loss for our person ID classification part is as follows [14]:
$\mathcal{L}^{I2T}_{id}&=\frac{1}{n}\sum\limits_{i=1}^{n}-log\big{(}\frac{e^{(\alpha(s_{y_{i},i}-\mathcal{M}_{2}))}}{e^{(\alpha(s_{y_{i},i}-\mathcal{M}_{2}))}+\sum\limits_{j\neq{y_{i}}}e^{(\alpha(s_{y_{j},i}))}}),\end{aligned}$ (8)
$s_{i,j}=\omega^{T}_{i}\hat{v}_{j},\quad\hat{v}_{i}=(v^{c}_{i})^{T}\bar{t}^{c}_{i}\cdot\bar{t}^{c}_{i},$ (9)
where $n$ is batch size, and $\omega$ represents the classification weight after normalization. $\hat{v}_{i}$ can be explained as the projection of image representation $v^{cls}_{i}$ onto the normalized text representation $\bar{t}^{c}_{i}$.
$\mathcal{L}^{T2I}_{match}$ and $\mathcal{L}^{T2I}_{id}$ are in the same form as above, but the change is focused on the text-to-image. Both matching and classification loss have identical decision boundaries. Equipped with our proposed cross-modal constraints, the Sew calibration loss can effectively reduce the gap between image and text feature distributions. Although the margin restrictions allow our model to learn better cross-modal representations, using a fixed margin $\mathcal{M}$ in all cases may not be flexible enough. A large margin constraint makes model learning difficult, while too small a margin does not impose a significant constraint. An adaptive margin guided by quality can be more effective. In TBPS, the texts come from the annotator’s descriptions of the images. Images are expected to have complete information, while texts have varying amounts of information. Thus, we adjust the margin value based on the quality of the text description. We argue that a less informative caption (i.e., a shorter caption) needs a smaller margin as a looser constraint. Based on this, we compute the adaptive margin for each image-text pair according to its text total tokens length $\mathcal{T}_{i}$:
$\mathcal{M}_{i}=\mathcal{M}_{min}+\frac{(\mathcal{M}_{max}-\mathcal{M}_{min})\cdot(\mathcal{T}_{i}-\mathcal{T}_{min})}{\mathcal{T}_{max}-\mathcal{T}_{min}},$ (10)
where $\mathcal{M}_{max}$ and $\mathcal{M}_{min}$ are upper and lower bounds of margins. $\mathcal{T}_{max}$ and $\mathcal{T}_{min}$ are bounds of the captions length, respectively. We set $\mathcal{T}_{max}$ and $\mathcal{T}_{min}$ according to the different dataset captions length distributions. After that, we utilize $\mathcal{M}_{i}$ to replace the fixed margin $\mathcal{M}$ above. We simply set $\mathcal{M}_{1}=\mathcal{M}_{2}=\mathcal{M}_{i}$.
III-B Masking Caption Modeling Loss
TBPS is a fine-grained cross-modal task, which means only caption-level discrimination is not enough. If the textual captions of two persons differ in a few words, a TBPS method can not retrieve a specific person without word-level discrimination. Although there are many works to establish word-level discrimination capacities, such methods are complex and limited, hindering performance boost. To solve this issue, we propose masking caption modeling to establish detailed image-text relationships. Furthermore, by utilizing MCM, our framework can perform more generic cross-modal learning.
Inspired by [40, 41], we add a masked prediction task on the text branch. Concretely, this loss is based on a cross-modal decoder architecture. We mask a portion of text tokens and replace these masked tokens with a learnable token vector. The text encoder inputs these text tokens and outputs the corresponding text features. The cross-modal decoder $f_{cd}$ learns to maximize the conditional likelihood of the masked text feature $t_{n}$ under latent image feature sequence $\{v_{i}\}$ and text feature sequence $\{t_{i}\}$:
$\mathcal{L}_{mcm}=-\sum_{n=1}^{N}\text{log }P_{\theta}(t_{n}|\{t_{i}\},\{v_{i}\}),$ (11)
where $N$ is the total masked token numbers in a caption.
Input: Image $I$ and text $T$; A batch of $n$ paired $\mathbb{G}=\{(I^{1},T^{1}),(I^{2},T^{2}),\cdots,(I^{n},T^{n})\}$
Output: Training loss $(\mathcal{L}_{sew},\mathcal{L}_{mcm})$  or  $(I_{cls},T_{cls})$
1 foreach $T^{i}$ in set $\mathbb{G}$ do
2      $T^{i}\leftarrow\text{mask}(T^{i})$;
3 end foreach
4for $i\leftarrow 1$ to $\mathbb{G}$ do
5       $(I^{i}_{cls},I^{i}_{1},\cdots,I^{i}_{c})\leftarrow ViT(T^{i})$;
6       $(T^{i}_{cls},T^{i}_{1},\cdots,T^{i}_{c})\leftarrow Bert(T^{i})$;
7       if Training Stage then
8             $(I^{i}_{cls},T^{i}_{cls})\leftarrow BatchNorm(I^{i}_{cls},T^{i}_{cls})$;
9             $I^{i}_{context},T^{i}_{context}\leftarrow(I^{i}_{1},\cdots,I^{i}_{c}),(T^{i}_{1},\cdots,T^{i}_{c})$;
10             $T^{i}_{context}\leftarrow Attn(T^{i}_{context})$;
11             $T^{i}_{cross}\leftarrow CrossAttn(T^{i}_{context},I^{i}_{context})$;
12             $\mathcal{L}_{sew}^{i}\leftarrow SewCalibration(I^{i}_{cls},T^{i}_{cls})$;
13             $\mathcal{L}_{mcm}^{i}\leftarrow MCM(T^{i}_{cross},T^{i})$;
14
15      else
16             $\{(I^{1}_{cls},T^{1}_{cls}),(I^{2}_{cls},T^{2}_{cls}),\cdots,(I^{i}_{cls},T^{i}_{cls})\}\leftarrow(I_{cls},T_{cls})$;
17
18       end if
19
20 end for
21if Training Stage then
22       return $(\mathcal{L}_{sew},\mathcal{L}_{mcm})$;
23
24else
25       return $(I_{cls},T_{cls})$;
26
27 end if
28
Algorithm 1 Traning Procedure of SCMM
As shown in Fig. 2, the cross-modal decoder part, the decoder $f_{cd}$ takes both unmasked text tokens and masked tokens in their original order as the input. The multi-head self-attention [17] first encodes the text features as $Q_{t},K_{t},V_{t}$, while the cross self-attention further improves the text features by taking into account the encoded image features as $K_{i},V_{i}$ for visual context. The final linear projection layer has the same number of output channels as the text vocabulary and computes the cross entropy loss between the reconstructed and original words only on masked text tokens. It should be noted that $f_{cd}$ is only used during training and not during inference.
III-C Total Loss
The total loss $\mathcal{L}$ we optimized in each iteration is as follows:
$\mathcal{L}_{sew}=\mathcal{L}_{match}^{I2T}+\mathcal{L}_{match}^{T2I}+\mathcal{L}_{id}^{I2T}+\mathcal{L}_{id}^{T2I},$
$\mathcal{L}=\lambda_{1}\mathcal{L}_{sew}+\lambda_{2}\mathcal{L}_{mcm},$ (12)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters to balance the different loss terms during training. The pseudocode of our framework pipeline is shown in Algorithm. 1.
IV Experiments
IV-A Datasets and Evaluation Metric
We evaluate SCMM on three benchmarks: CUHK-PEDES [10], ICFG-PEDES [19], and RSTPReid [26]. CUHK-PEDES is the first large-scale benchmark for text-based person search tasks. This dataset contains 40,206 images of 13,003 person IDs collected from five person re-identification datasets. Each image has two different textual captions with an extensive vocabulary, and the average sentence length is 23.5. The testing set comprises 3,074 images and 6,148 descriptions of 1,000 persons. ICFG-PEDES contains 54,522 images of 4,102 persons. For each image, the corresponding description has an average length of 37 words. The testing set consists of 19,848 image-text pairs of 1,000 persons. RSTPReID contains 20,505 images of 4,101 persons, with each pedestrian having five images. It is divided into a training set with 3,701 persons, a validation set with 200 persons, and a testing set with 200 persons. For our evaluation metric, we report the Rank $k$ ($k$=1, 5, 10) text-to-image accuracy, which is commonly used in previous works to evaluate text-based person search. Given a textual description as the query, if the top-$k$ retrieved images contain any person corresponding to the query, we consider it a successful person search.
IV-B Implementation Details
In our visual-textual dual-encoder, we extract visual representations using the ViT-Base pre-trained on ImageNet [42]. The images are resized to 224 $\times$ 224 pixels. For textual representations, we use the BERT-Base-Uncased model pre-trained on the Toronto Book Corpus and Wikipedia. The representation dimension is set to 768, and the feature sequence lengths are set to 197 and 100, respectively.
During the training phase, we use a batch size of 64 and train for 60 epochs. We use Adam optimizer with an initial learning rate of 0.001. To augment our data, we apply a random horizontal flipping operation, and we use a mask ratio of 0.1 for randomly masking text tokens. The minimum and maximum textual information length boundaries $\mathcal{T}_{min}$ and $\mathcal{T}_{max}$ are set to 20-60, 25-65, and 22-60 according to the caption length distributions of CUHK-PEDES, ICFG-PEDES, and RSTPReID, respectively. The bounds of the upper and lower margin $\mathcal{M}$ are set to 0.4 and 0.6, and the scale parameter $\alpha$ is set to 32. For each loss in the total loss function, the balance factors $\lambda_{1}$ and $\lambda_{2}$ are set equal to 1. During the testing phase, we apply a re-ranking post-processing approach to improve search performance following NAFS [28]. We conduct the experiments on four NVIDIA 2080Ti GPUs using PyTorch.
TABLE I: Comparison with SOTA methods on the CUHK-PEDES dataset. Rank1, Rank5, and Rank10 accuracies (%) are reported.
Method Rank1 Rank5 Rank10
GNA-RNN [10] 19.05 - 53.64
CMPM/C [14] 49.37 - 79.27
Dual-path [12] 44.40 66.26 75.07
ViTAA [18] 55.97 75.84 83.52
VP Net [27] 58.83 81.25 86.72
HGA Net [43] 59.00 79.49 86.62
SUM [44] 59.22 80.35 87.60
CLIP [29] 60.67 81.99 88.87
DSSL [26] 59.98 80.41 87.56
DSSL+R [26] 62.33 82.11 88.01
MGEL [22] 60.27 80.01 86.74
SSAN [19] 61.37 80.15 86.73
TestReID [45] 64.08 81.73 88.19
TestReID+R [45] 64.40 81.27 87.96
ACSA [25] 63.56 81.40 87.70
ACSA+R [25] 68.67 85.61 90.66
IVT [39] 64.00 82.72 88.95
SAFA Net [21] 64.13 82.62 88.40
TIPCB [20] 64.26 83.19 89.10
CAIBC [46] 64.43 82.87 88.37
AXM Net [47] 64.44 80.52 86.77
CFine [24] 65.07 83.01 89.00
CFine+C [24] 69.57 85.93 91.15
TBPS-CLIP [48] 65.72 84.62 90.96
ISANet [8] 63.92 82.15 87.69
SCMM 67.71 84.57 89.44
SCMM+R 71.09 86.78 91.23
SCMM+C 69.61 86.01 90.90
SCMM+R+C 73.81 88.89 92.77
  • Bold number represents the best score. R and C denote the re-ranking post-processing operations and image encoder pre-trained on CLIP model.
TABLE II: Quantitative results on the ICFG-PEDES dataset.
Method Rank1 Rank5 Rank10
CMPM/C [14] 43.51 65.44 74.26
Dual-path [12] 38.99 59.44 68.41
ViTAA [18] 50.98 68.79 75.78
SSAN [19] 54.23 72.63 79.53
CLIP [29] 53.96 73.69 80.43
TIPCB [20] 54.96 74.72 81.89
IVT [23] 56.04 73.60 80.22
CFine [24] 55.69 72.72 79.46
CFine+C [24] 60.83 76.55 82.42
IRRA [49] 63.46 80.25 85.82
TP-TPS [50] 60.64 75.97 81.76
ISANet [8] 57.73 75.42 81.72
SCMM 60.20 75.97 81.78
SCMM+R 62.17 85.74 89.67
SCMM+C 62.29 77.15 82.52
SCMM+R+C 64.25 86.95 90.70
IV-C Comparison with State-of-the-art Methods
IV-C1 Results on CUHK-PEDES.
Table I compares our framework and previous methods on CUHK-PEDES. It can be observed that our proposed SCMM can outperform all previous methods by a large margin. Compared to the state-of-the-art work CFine [24], SCMM achieves 67.71%(+2.64%), 84.57%(+1.56%) and 89.44%(+0.44%) on Rank1, Rank5 and Rank10 without re-ranking. The state-of-the-art results on the CUHK-PEDES show the effectiveness of SCMM. With the help of re-ranking, SCMM shows an incremental boost to get a 71.09% Rank1 score. Furthermore, with a better image encoder pre-training on CLIP, SCMM achieves 73.81%, 88.89% , and 92.77% on three metrics, respectively. These consistent improvements show the scalability of SCMM across better pre-trained models and extra post-processing operations.
IV-C2 Results on ICFG-PEDES and RSTPReid.
We also utilized other benchmarks to validate SCMM’s performance and generalization. The ICFG-PEDES and RSTPReid datasets are more challenging compared to CUHK-PEDES, and our method significantly outperformed all state-of-the-art methods on these two datasets by a large margin, as reported in Tables II and III. Compared with the state-of-the-art results [24] on ICFG-PEDES, SCMM achieved significant improvements, with scores of 60.20%(+4.51%), 75.97%(+3.25%), and 81.78%(+2.32%) on Rank1, Rank5, and Rank10, respectively. On the RSTPReid dataset, we achieved scores of 50.75% (+4.90%), 74.20%(+3.90%), and 81.70%(+3.30%) on the three metrics. With re-ranking post-processing, we were able to achieve scores of 62.17% and 55.35% on Rank1 for the two benchmarks, respectively.
We note that re-ranking also brings a significant boost, as our model learns clear and compact cross-modal key information patterns, and re-ranking can retrieve the correct feature neighbors more effectively. Moreover, by utilizing the CLIP pre-trained image model as an encoder, we achieved even better results, with scores of 64.25%, 86.95%, and 90.70% on ICFG-PEDES, and scores of 57.35%, 77.50%, and 85.50% on RSTPReid. These two challenging benchmark results demonstrate the robustness and scalability of SCMM.
TABLE III: Quantitative results on the RSTPReid dataset.
Method Rank1 Rank5 Rank10
CLIP [29] 50.10 76.10 84.95
DSSL [26] 32.43 55.08 63.19
SUM [44] 41.38 67.48 76.48
SSAN [19] 43.50 67.80 77.15
IVT [23] 46.70 70.00 78.80
ACSA [25] 48.40 71.85 81.45
CFine [24] 45.85 70.30 78.40
CFine+C [24] 50.55 72.50 81.60
TP-TPS [50] 50.65 72.45 81.20
SCMM 50.75 74.20 81.70
SCMM+R 55.35 77.30 84.25
SCMM+C 51.95 73.50 82.45
SCMM+R+C 57.35 77.50 85.50
Figure 4: Comparison of singular values for image-text embedding features across CUHK-PEDES (a-b) and ICFG-PEDES (c-d) datasets. (a) and (c) depict the distribution of singular values, where a smaller inter-line gap indicates a closer cross-modal distribution. (b) and (d) present the logarithmic difference in singular values between the baseline and our approach for CUHK-PEDES and ICFG-PEDES, respectively.
IV-D Ablation Studies
IV-D1 Analysis of Model Components
To fully validate the performance of the different components in SCMM, we demonstrate the contributions of each part on CUHK-PEDES, as shown in Table IV. Model 1 and Model 2 show the results using ResNet [51] and ViT [39] as the image encoder, respectively, without the Sew Calibration and MCM losses. We utilize Model 2 as our baseline, and CMPM and CMPC [14] are used as loss functions in the baseline experiment. First, we compare ResNet-50 and ViT-Base as the image encoder and observe a performance improvement of 1.90%, 1.03%, and 0.69% on Rank1, Rank5, and Rank10, respectively. Based on this, we adopt ViT as the image encoder in the following experiments. We then validate the Sew calibration and MCM losses compared to the baseline. From Models 2-4, we observe that Sew calibration loss brings marginal improvement, and a common distribution can provide a better basis for optimizing the embedding space for fine-grained cross-modal recognition. Moreover, Model 3 represents the Sew calibration loss with the fixed margin 0.5, while Model 4 represents the Sew calibration loss with adaptive margins. We observe that Model 4 achieves better results than a fixed manual margin, no matter how we tune the margin value, which demonstrates the effectiveness of adaptive constraints. We also find that MCM gives another substantial improvement to the baseline from Model 2 to Model 5, indicating that the text-image detail mining capability is critical. Notably, when the two components are used jointly in Model 6, SCMM continues to improve and outperform the baseline by 4.57%, 2.99%, and 2.37% on the three metrics, respectively. This demonstrates that obtaining well-distributed image-text features in the common embedding space is essential for reducing the cross-modal gap. The consistent improvement in each component of the method demonstrates our effectiveness.
TABLE IV: Performance comparisons of different components.
ID Res ViT Sew-F Sew-A MCM Rank1 Rank5 Rank10
1 60.39 80.22 86.53
2 62.29 81.25 87.22
3 65.41 82.56 88.01
4 66.02 83.33 88.61
5 65.16 81.95 87.93
6 67.71 84.57 89.44
  • Res and ViT denote ResNet and Vision Transformer as image encoders,
  • respectively. Sew-F and Sew-A refer to the fixed and adaptive margin
  • versions of the Sew calibration loss, respectively. Model 2 is our baseline.
IV-D2 Impact of Sew Calibration Loss for Reducing Cross-modal Gap
Benefiting from the Sew calibration loss, which reduces the cross-modal gap, the learned representation distributions are closer than with the basic loss. To demonstrate this, we illustrate a comparison of singular value decomposition on cross-modal representations, inspired by [52]. Specifically, Fig. 4(a) and (c) present the baseline and SCMM’s singular value decomposition for the text and image modalities. We computed the distance between these two modalities at specific singular values and reflected them in Fig. 4(b) and (d). Intuitively, the smaller the distance between different modality features, the closer the distribution learned by the model. We find that the Sew calibration loss performs better representation distributions in the common embedding space. It ensures closer cross-modal representation distributions and a smaller inter-modal gap than the baseline.
TABLE V: MCM components on the CUHK-PEDES dataset.
Mask Attention Caption Modeling Rank1 Rank5 Rank10
63.65 81.84 87.51
63.83 80.87 86.35
64.28 81.97 87.75
65.16 81.95 87.93
IV-D3 Impact of Masking Caption Modeling
The masking caption modeling operation in fine-grained cross-modal interaction is critical in our framework. It comprises tokens masking, attention module, and masked tokens prediction. As shown in Table V, we explore the effectiveness of MCM components on the CUHK-PEDES. First, only masking on the input text tokens without a reconstruction task behaves like a random erase text augmentation. This augmentation can already bring +1.36% marginal improvement and reach 63.65% on Rank1. It shows that the details provided by word tokens are helpful for fine-grained recognition.
Next, if we only utilize attention mechanism to enhance cross-modal representations without the mask caption modeling, it can achieve 63.83% on Rank1. We explain this improvement as the cross-attention brings details from sequence tokens to the CLS token. Meanwhile, image features also contribute a lot to this process.
On the other hand, we also design an experiment of mask caption modeling without the cross attention. The decoder $f_{cd}$ directly predicts all masked tokens from dual-encoder outputs. In this experiment, we observe 64.28% on Rank1. It proves the reconstruction tasks in the decoder help guide the text encoder to learn richer and more refined representations. We also notice that with all three components, the MCM achieves 65.16% on Rank1. With the help of cross attention from image features, the CLS token ensembles all those rich information.
Figure 5: Visualization of attention maps from baseline and SCMM on the CUHK-PEDES. We present the total caption-level results and the fine-grained word-level results, respectively. (1) Baseline caption-level results, (2) Our caption-level results, (3) & (4) Our word-level results. Best viewed in color.
TABLE VI: Performance analysis for cross-domain validation.
CUHK $\Rightarrow$ ICFG Rank1 Rank5 Rank10
Baseline 30.79 49.10 57.47
Baseline+MCM 31.13 49.52 58.07
ICFG $\Rightarrow$ CUHK Rank1 Rank5 Rank10
Baseline 22.19 40.55 50.52
Baseline+MCM 23.78 42.41 52.11
CUHK $\Rightarrow$ RSTP Rank1 Rank5 Rank10
Baseline 37.40 63.40 74.35
Baseline+MCM 39.25 63.95 74.55
RSTP $\Rightarrow$ CUHK Rank1 Rank5 Rank10
Baseline 10.02 22.95 31.04
Baseline+MCM 10.69 25.20 33.93
IV-D4 Analysis of Generalisability Validation
To better understand the contributions of MCM in our framework, we conduct a domain generalization analysis on the three benchmarks to demonstrate our generalization performance. In Table VI, CUHK $\Rightarrow$ ICFG and CUHK $\Rightarrow$ RSTP indicate using the model trained on the CUHK-PEDES dataset to infer their test sets and vice versa. Compared to the baseline, we can observe a performance improvement in the three metrics for ICFG-PEDES and RSTPReid. The fine-grained features mined by MCM are more resistant to overfitting. The improvement in domain generalization shows the capability of our framework in learning generic cross-modal information, which is essential to solve the fine-grained modal heterogeneity.
IV-D5 Effect of Margin Parameters
In subsection  III-A, we introduce the adaptive margins to constrain cross-modal learning using Sew Calibration loss with constraints. We find that compared to the manually fixed parameters (SEW-f), our adaptive margins (SEW-a) produce better results as shown in Fig. 6, where we investigate the effect of manual margin in different settings on the CUHK-PEDES [10] on Rank1 accuracy. The best performance of 65.41% is achieved when $\mathcal{M}=0.5$, while the worst model performance of 62.29% is observed when $\mathcal{M}=0$. We note that the performance of the model first improves as the margin increases. This is due to our bi-directional margin compressing the same pedestrian target compactly. However, we find that the model performance does not continue to improve as the margin increases, and we observe an inflection point around margin 0.5. When the margin is too large, the tight constraints make the model too hard to train.
Therefore, it is critical to utilize quality-guided adaptive margins according to the textual information. In this regard, we set the upper and lower margin bounds ($\mathcal{M}_{min}=0.4$ and $\mathcal{M}_{max}=0.6$). In this range, our Sew Calibration loss with adaptive constraints obtains the best performance, which a manually fixed margin cannot achieve, regardless of how the margin value is tuned.
IV-D6 Effect of Mask Ratio
Inspired by Masked Language Modeling [37], we propose the MCM loss which leverages a masked captions prediction task to establish detailed and generic relationships between textual and visual parts. Our method requires setting the mask ratio in the MCM. To investigate the effect of the mask ratio on model performance and generalization learning ability, we conducted ablation experiments, and the results are shown in Table VII. When the mask ratio is set to 0, which is equivalent to the baseline, the performance is the worst. Compared to the baseline, our method with a 0.1 mask ratio achieves the best results, with scores of 65.16%, 81.95%, and 87.93% on Rank1, Rank5, and Rank10, respectively.
TABLE VII: Ablation results of mask ratio on the CUHK-PEDES.
Baseline+MCM Rank1 Rank5 Rank10
Mask = 0 62.29 81.25 87.22
Mask = 0.1 65.16 81.95 87.93
Mask = 0.3 64.97 81.38 87.71
Mask = 0.5 63.78 81.11 87.43
The experiments are conducted on the CUHK-PEDES dataset using the Baseline+MCM method. As shown in these results, we find that the model achieves the best performance when the mask is set to 0.1, and then the model performance gradually decreases as the mask ratio increases. This is because the information provided in the annotated caption text is very accurate and free of redundant information. Therefore, as the mask ratio expands, more and more key information may be obscured, causing the model to fail to learn enough semantic information to complete the construction of the image-text relationship.
Figure 6: Effect of the manual fixed margin parameters setting of our Sew Calibration loss in terms of Rank1 accuracy on the CUHK-PEDES.
IV-E Qualitative Analysis
IV-E1 Visualization of Attention Map
We visualize attention maps on the CUHK-PEDES test dataset to demonstrate model capability in learning image-text correspondence. As shown in Fig. 5, compared to the baseline, we can observe that the visualization results obtained by SCMM are more apparent and refined. We conduct the word-level visualization to validate further the ability to perform fine-grained interaction. We select several keywords in the caption description, e.g., bag, shoes as items, colors, and clothing as modifiers. For example, the visualization results of IDs 8491 and 10648 do not focus on useful detailed information. The key messages in the two images are a man riding bike with a backpack and a woman wearing a sleeveless white dress. We can observe that SCMM successfully captures the key detail information compared to the baseline. From the word-level results, we can also observe that SCMM learns the critical parts in the cross-modal correspondence.
Figure 7: Presentation of cross-modal representation distributions on CUHK-PEDES test dataset. Different colors correspond to the different target ID. Best viewed in color.
IV-E2 Visualization of Image-text Feature Distribution
SCMM is capable of learning better image-text representation distributions with fine-grained image-text interaction and adaptive constraints to reduce the cross-modal gap. To demonstrate this, we randomly selected some pedestrians and extracted their image-text global features, then mapped them to two dimensions for visualization in Fig. 7(a) and (b). For instance, we can take IDs 5532 and 11800 as examples. We can observe that the image-text features of the same ID distribution are compressed more compactly, and the boundaries between them are more apparent than the baseline results. This demonstrates the effectiveness of SCMM in mitigating the cross-modal gap.
V Conclusion
In this work, we introduce SCMM, a novel method for TBPS that addresses the critical need for precise and efficient person retrieval in IoT environments. SCMM enhances cross-modal representation learning to reduce the inter-modal gap and discern fine-grained image-text correspondences. This is achieved through a dual-pronged strategy employing Sew calibration and MCM losses, which together refine feature alignment and enrich cross-modal correspondences. The method’s streamlined architecture ensures high-speed inference, essential for real-time applications in IoT settings. For instance, SCMM can be integrated with edge computing units in surveillance cameras to preprocess video streams and extract textual descriptions of individuals, facilitating efficient person search across the network. This integration is crucial for applications in smart cities and industrial environments, where timely and accurate person identification is vital for maintaining security and operational efficiency. Extensive experiments on three popular TBPS benchmarks demonstrate the effectiveness and superiority of SCMM. Future work will explore enhancing SCMM’s generalizability to multi-modal inputs, investigating its potential in multi-view TBPS scenarios, and addressing challenges in domain adaptation and few-shot learning, thereby advancing TBPS towards more practical and versatile applications in smart surveillance and multimedia ecosystems.