### v7

# SCMM: Calibrating Cross-modal Representations for Text-Based Person Search

Abstract
Text-Based Person Search (TBPS) represents a critical engineering application of artificial intelligence (AI) in intelligent surveillance systems where automated person identification using natural language descriptions is essential. Current AI methods face significant limitations in achieving well-aligned cross-modal representations and capturing fine-grained correspondences essential for discriminating similar individuals. The fundamental challenge lies in developing intelligent systems that effectively process heterogeneous data sources where visual and textual features exist in different semantic spaces, creating substantial inter-modal gaps. To address these challenges, we propose Sew Calibration and Masked Modeling (SCMM), a novel AI framework that calibrates cross-modal representations through strategic techniques. SCMM introduces two innovative mechanisms: a sew calibration loss that dynamically aligns image-text features based on textual information quality using adaptive margin learning, and a masked caption modeling loss that establishes detailed cross-modal correspondences through transformer-based masked prediction tasks. The sew calibration component implements quality-aware adaptive constraints, enabling fusion strategies that accommodate varying information richness across modalities. While the masked modeling mechanism facilitates fine-grained feature learning by discovering detailed relationships between visual elements and textual descriptions through cross-modal attention. Our streamlined dual-encoder architecture maintains computational efficiency suitable for real-time applications while achieving superior performance through complementary alignment and correspondence learning strategies. Extensive experiments across three benchmark datasets demonstrate SCMM’s effectiveness, achieving superior Rank1 accuracy on three challenging benchmarks. The framework validates the practical applicability of quality-aware adaptive constraints and fine-grained correspondence modeling for real-world AI engineering applications.
keywords:
Text-Based Person Search, Cross-modal Learning, Person Re-identification, Metric Learning, Engineering Applications, Artificial Intelligence
journal: Engineering Applications of Artificial Intelligence
[…]
[label3]organization=Department of ECE, The University of British Columbia, city=Vancouver, postcode=V6T 1Z4, country=Canada
[…]
[label1]organization=Academy for Engineering &\&& Technology, Fudan University, city=Shanghai, postcode=200433, country=China
[…]
[label4]organization=Department of Computer Science, University of Toronto, city=Toronto, postcode=M5S 1A1, country=Canada
[…]
[label2]organization=MEGVII Technology, city=Beijing, postcode=100096, country=China
[…]
[label7]organization=Faculty of Data Science, City University of Macau, city=Macau, postcode=999078, country=China
1 Introduction
Artificial intelligence (AI) applications in engineering have revolutionized modern computer vision and machine learning systems, driven by the urgent need to develop intelligent solutions that can process and integrate heterogeneous data sources for enhanced understanding and comprehensive analysis capabilities [1, 2]. Person re-identification (Re-ID) exemplifies a critical engineering application of AI, requiring sophisticated machine learning algorithms to process visual features across non-overlapping camera views for identifying target individuals in crowded scenes and enhancing security measures in smart city infrastructure [3]. However, traditional Re-ID approaches face significant operational limitations in real-world engineering scenarios when query images are unavailable, necessitating the development of alternative AI-driven solutions [4]. Text-Based Person Search (TBPS) addresses this critical engineering challenge by employing advanced artificial intelligence techniques to process textual descriptions alongside visual galleries, representing a significant advancement in practical AI applications for surveillance and security systems [5]. The implementation of multi-modal processing in TBPS enables the effective integration of rich semantic information from natural language descriptions with visual features from surveillance imagery [6], thereby providing intelligent solutions for Re-ID in scenarios where visual queries are unavailable or insufficient. Consequently, the paradigm shift from traditional visual-only approaches to sophisticated AI-powered cross-modal techniques addresses fundamental limitations in real-world deployment while opening new possibilities for intelligent surveillance and security applications [7].
The effectiveness of AI-powered multi-modal systems depends critically on developing machine learning algorithms that can bridge the semantic gap between heterogeneous modalities while preserving discriminative information essential for accurate person identification [8]. Engineering challenges in TBPS stem from the fundamental differences in feature representations between textual and visual domains, where semantic concepts must be aligned across disparate embedding spaces using advanced AI techniques [9]. State-of-the-art AI architectures have emerged to address these engineering challenges, incorporating transformer-based attention mechanisms and deep learning models for cross-modal interactions to enhance intelligent processing capabilities [10]. Contemporary research in AI engineering applications emphasizes the development of sophisticated alignment strategies that can effectively combine complementary information sources while maintaining computational efficiency suitable for practical deployment in real-world systems [11]. The integration of these advanced AI techniques has proven essential for engineering applications requiring precise person identification with natural language queries, particularly in scenarios where traditional visual-only approaches prove insufficient for accurate target discrimination in intelligent surveillance systems [12].
Figure 1: Illustration of the motivation behind our method. In cross-modal tasks, a compact and well-aligned image-text feature distribution in the shared embedding space is crucial for bridging the inter-modal gap. Additionally, capturing fine-grained image-text correspondences is equally vital to distinguish between similar individuals for text-based person searching.
Compared to conventional Re-ID, the main objective of TBPS is to learn fine-grained cross-modal representations between visual and textual modalities [13]. As shown in Fig. 1, cross-modal representations have two key characteristics that contribute to effectively searching for a person from images: closely aligned cross-modal representations and fine-grained information correspondence. First, closely aligned cross-modal representations can effectively reduce the inter-modal gap, making it easier to locate specific persons using textual captions. Second, detailed correspondence is absolutely essential for TBPS, as fine-grained cross-modal information is often necessary to discriminate between similar persons [14].
In the TBPS domain, numerous approaches have been proposed from two distinct perspectives. First, some methods [15, 16] employ only dual encoders and align the two modal representations using symmetric loss functions. Although computationally efficient, these methods suffer from limited feature alignment capabilities. Alternative approaches [17] utilize multi-modal models with transformer-based [18] cross-attention mechanisms to improve cross-modal feature alignment and interaction. However, such methods often require high computational costs due to the exhaustive fusion of all possible image-text pairs during inference, hindering their efficiency in real-time applications. Second, to obtain robust fine-grained features, recent TBPS works have designed multi-level [19, 20, 21, 22], multi-granularity [23, 24] matching strategies, and specialized attention modules [8, 25, 26]. While these methods provide fine-grained features, they suffer from complex model architectures and expensive computations. Furthermore, the fine-grained features they produce are often limited and hinder significant performance improvements.
To address these persistent challenges, we propose a simple yet effective method called SCMM for calibrating cross-modal representations in TBPS. SCMM employs only a dual-encoder architecture, making it computationally efficient without requiring complex interaction modules or additional multi-level branches, which enables high-speed inference suitable for practical applications. Additionally, we propose two novel training losses to effectively calibrate cross-modal representations. The first is a sew calibration loss, which leverages the quality of text descriptions as guidance to align features between textual and visual modalities while pushing negative sample pairs apart and pulling positive sample pairs together across modalities. Furthermore, we propose a Masked Caption Modeling (MCM) loss to obtain more fine-grained and generalizable correspondence. Building upon the principle of masked caption prediction, the loss employs a masked caption prediction task to establish detailed relationships between textual and visual components. The operation is implemented through a cross-modal decoder that is discarded during inference, avoiding additional computational costs. To comprehensively demonstrate SCMM’s effectiveness, we evaluated its performance on three popular benchmarks, where our model surpasses previous state-of-the-art (SOTA) methods and demonstrates impressive performance. Moreover, we conducted extensive experiments to systematically validate each component of our method. Overall, our main contributions can be summarized as follows:
  • 1.
    We propose an efficient and scalable framework for text-based person search, named SCMM, that calibrates cross-modal representations using a dual-encoder architecture and an auxiliary cross-modal decoder. The framework’s design effectively resolves the trade-off between the high computational cost of complex interaction models and the limited alignment capability of simpler methods, achieving superior performance with high inference speed suitable for real-world AI engineering applications.
  • 2.
    We introduce a novel sew calibration loss, which uniquely incorporates a quality-guided adaptive margin mechanism. By dynamically adjusting alignment constraints based on the information density of textual descriptions, our loss function more effectively bridges the inter-modal gap, pushing visual and textual feature distributions into a more compact and well-aligned shared embedding space.
  • 3.
    We propose an innovative MCM loss that establishes detailed cross-modal correspondences through masked prediction tasks and transformer-based cross-modal attention mechanisms. Our method facilitates fine-grained information integration by learning word-level relationships between visual elements and textual descriptions, enabling the framework to distinguish between similar individuals with subtle differences while maintaining computational efficiency through training-only decoder utilization.
  • 4.
    We conduct extensive experiments on three challenging benchmarks: CUHK-PEDES, ICFG-PEDES, and RSTPReID. SCMM demonstrates superiority by achieving new state-of-the-art Rank1 accuracies of $73.81\%$, $64.25\%$, and $57.35\%$ respectively. The establishment of new performance benchmarks not only validates the effectiveness of our fusion strategies for text-based person search, but also provides valuable baselines for future research in cross-modal AI applications.
The structure of the remainder of this paper is outlined as follows. Sec. 2 reviews the related work on text-based person search, metric learning, and masked language modeling. Sec. 3 introduces the proposed SCMM method, including the sew calibration loss, masked caption modeling loss, and the overall framework. Sec. 4 presents the experimental setup, datasets, implementation details, and comprehensive experimental results, including comparison with state-of-the-art methods, ablation studies on different components, and qualitative analysis with visualization of attention maps and feature distributions. Sec. 5 concludes the paper with a summary of contributions and future research.
2 Related Work
2.1 Text-Based Person Search
Text-based person search was first introduced by [8], which identifies person images in a gallery using only textual queries. Fundamentally, the task represents a critical application of multi-modal fusion, where the integration of heterogeneous data sources—visual and textual modalities—becomes essential for enhanced understanding and analysis capabilities. Early works utilized pre-processing methods to obtain external cues such as person segmentation [19] and human body landmarks [27]. The evolution towards sophisticated fusion architectures has become crucial, as multi-view fusion combining complementary information from multiple sources shows significant promise for enhancing prediction accuracy and robustness in cross-modal tasks [10].
In recent years, end-to-end frameworks based on attention mechanisms [22, 24, 26] have become prevalent, addressing the fundamental challenge of semantic mismatch between features of different modalities. Cross-modal attention is critical for performing image-text interaction, where advanced fusion strategies must consider the unique attributes of cross-modal matching tasks [3]. Existing methods can be broadly classified into attention-explicit and attention-implicit approaches. Attention-explicit methods [2, 26] design specific attention modules according to multi-granularity and multi-level strategies, implementing early fusion techniques that combine features at the input level. For example, NAFS [14] conducts cross-modal alignments over full-scale features with a contextual non-local attention module, while CFine [25] utilizes cross-attention for multi-grained global feature learning, achieving impressive results through knowledge transfer from the CLIP [28] model.
In contrast, attention-implicit methods [22, 24] utilize transformer-based models with shared parameters to align cross-modal semantics implicitly, representing late fusion approaches that combine decisions from individual modality-specific models [10]. However, compared to performing in-depth cross-attention with task-driven approaches, existing methods face challenges in addressing information asymmetry problems and do not implement sufficient cross-interaction with generalized performance, highlighting the need for more sophisticated fusion techniques that enhance feature integration and establish deeper connections between global structure and local details [3].
2.2 Metric Learning
Initially, metric learning used $L_{2}$ distance as the metric, with the goal is to minimize the $L_{2}$ distance between samples of the same class. Some $L_{2}$-based metric learning methods include Siamese Networks [29] and Triplet Networks [30]. With the development of deep learning, researchers started using softmax-based loss functions in metric learning to learn a more discriminative distance metric. Additionally, increasing the margin between classes is an intuitive approach to learn a better metric space [3]. L-Softmax [31] introduced the concept of margin on softmax function for the first time. The widely used CosFace [32] proposed large-margin cosine loss to learn highly discriminative deep features for face recognition. Circle loss [33] proposed a simple loss based on a unified loss for metric learning and classification. Recently, some works introduced adaptive margin into marginal loss [2]. They usually learn image quality implicitly and adjust the margin accordingly. In single-modal representation learning, they usually give large margins to high-quality samples for hard mining. Compared to them, we try to solve a cross-modal matching problem where samples from two modalities have different information volume. We give greater tolerance to less informative samples in TBPS.
However, existing metric learning approaches face significant limitations when applied to cross-modal scenarios. First, traditional methods primarily focus on single-modal optimization and fail to adequately address the heterogeneity gap between different modalities [34]. Second, fixed margin constraints cannot adaptively handle the varying information density across modalities, leading to suboptimal alignment performance. Finally, most current approaches lack fine-grained correspondence modeling capabilities essential for distinguishing between similar cross-modal samples [4].
2.3 Masked Language Modeling
MLM is a highly effective method for pre-training language models [35] by randomly selecting a certain percentage of words from the input sentence and then predicting the masked words based on the context of other words. Many cross-modal pre-training models have utilized MLM in their methods. For example, the work in [17] combines MLM with contrastive loss in the framework, which achieved impressive performance in their cross-modal tasks. The success of MLM in BERT [36] has proven its ability to adapt well to various downstream tasks, leading to generalized performance. In our fine-grained framework, the task-driven decoder utilizing masked caption modeling can facilitate generic cross-modal learning.
However, existing MLM-based cross-modal approaches face several limitations when applied to fine-grained person search tasks. First, traditional MLM methods primarily focus on intra-modal language understanding and lack explicit cross-modal interaction mechanisms, limiting their ability to establish detailed visual-textual correspondences [6]. Second, most existing approaches apply MLM uniformly without considering the varying information density and quality of textual descriptions, leading to suboptimal learning efficiency. Finally, current methods often require complex multi-stage training procedures or heavy computational overhead during inference, hindering their practical deployment in real-time applications [23].
3 Proposed Methods
Formally, given a set of images with corresponding captions, denoted as $X=\{(I_{i},T_{i})\}^{N}_{i=1}$, where each image $I_{i}$ and its description text $T_{i}$ is associated with a person ID $y_{i}$. The text-based person search task aims to retrieve the most relevant Rank $k$ (e.g., $k=1,5,10$) person images efficiently from a large-scale gallery using textual queries. As an engineering application of AI, it requires developing machine learning algorithms that can bridge the semantic gap between visual and textual modalities while maintaining discriminative capabilities for accurate person identification in real-world scenarios. The fundamental engineering challenge lies in designing AI systems that can align heterogeneous feature representations across modalities, where visual features capture spatial and appearance characteristics while textual descriptions encode semantic attributes and contextual information. Effective strategies must address both global alignment between modalities and fine-grained correspondences that enable discrimination between similar individuals in practical applications.
Figure 2: Overview of our proposed SCMM. The framework consists of a dual-encoder for extracting image-text features and calibrating cross-modal representations with the sew calibration loss. We also include a decoder for performing cross-modal interaction with the task-driven Mask Caption Modeling. At the inference stage, we only utilize the classification (CLS) tokens from the dual-encoder to implement similarity search.
To address these engineering challenges in AI applications, we propose SCMM, a streamlined yet effective machine learning framework as illustrated in Fig. 2. Our approach implements a dual-encoder architecture using state-of-the-art AI models: ViT [37] and BERT [36] as backbone networks for visual and textual feature extraction respectively. The image encoder processes image patches from $I_{i}$ along with a vision classification (CLS) token, outputting an image feature sequence $\{v_{i}\}$ and a vision CLS token embedding $v_{i}^{c}$. Similarly, the text encoder generates a text feature sequence $\{t_{i}\}$ and a text CLS token embedding $t_{i}^{c}$ for caption $T_{i}$, following established AI engineering practices [14, 25]. Our framework integrates two complementary fusion mechanisms to calibrate cross-modal representations: the sew calibration loss with adaptive margin constraints for global alignment, and the masked caption modeling loss for establishing detailed cross-modal correspondences. The following subsections elaborate on each component. We first introduce the sew calibration loss that dynamically aligns cross-modal features based on textual information quality using machine learning techniques. Subsequently, we present the masked caption modeling mechanism that facilitates fine-grained information integration through transformer-based cross-modal attention and masked prediction tasks. The total loss function combines these complementary AI objectives to jointly optimize global cross-modal alignment and fine-grained correspondence learning, enabling the framework to achieve superior performance while maintaining computational efficiency suitable for real-world engineering applications through the streamlined dual-encoder architecture.
3.1 Sew Calibration Loss with Constraints
In the TBPS task, closely aligning cross-modal representations is crucial for effectively finding specific persons with textual captions. To address the heterogeneity between modalities, we propose a sew calibration loss that pushes each modality to a common space. As shown in Fig. 3(a), in single modality settings, a triplet distance constraint is widely utilized when embeddings are from a single modality distribution. In intra-class samples, the embeddings are pulled together, while inter-class samples are pushed away. In the cross-modal setting, we expect to impose constraints from both sides of the two embedding distributions. For example, in one-direction retrieval, we first need to align the features of a ”perfect pair,” i.e., $(I_{i},T_{i})$. The shortest distance from the image embedding ${v_{i}}$ to the text embedding should be ${t_{i}}$. Consequently, it is a perfect matching pair in person re-identification, and no other text feature will have a shorter distance (Eq. 1). We set the perfect pair as an image-text anchor $(A_{img},A_{txt})$. Next, we impose another constraint between the image anchor $A_{img}$ and its corresponding positive text samples $P_{txt}(\mathds{1}(y_{i}=y_{j},i\neq j))$ and negative text samples $N_{txt}$ (Eq. 2). For the other direction, we put symmetric constraints on $A_{txt}$ and $P_{img}$, $N_{img}$. Fig. 3(b) shows the proposed constraints for cross modality. Forces from both sides act like a seam to pull the two distributions together.
(a) Single-modal constraints
(b) Cross-modal constraints
Figure 3: Illustration of sew calibration loss. The constraints are different between single-modal and cross-modal matching. $(A_{img},A_{txt})$ denotes anchors in image-text feature distribution, while $(P_{img},P_{txt})$ and $(N_{img},N_{txt})$ denote positive and negative sample pairs, respectively. The sew calibration loss pushes negative sample pairs and pulls positive sample pairs, stitching cross-modal key information like a seam.
Take the image side as an example, the $L_{2}$ distance constraints are shown as follows:
$D(A_{img},A_{txt})+\mathcal{M}_{1}<D(A_{img},P_{txt}),$ (1)
$D(A_{img},P_{txt})+\mathcal{M}_{2}<D(A_{img},N_{txt}),$ (2)
where $D$ denotes $L_{2}$ distance. $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are two margins, $\mathcal{M}_{1}<\mathcal{M}_{2}$. With a bi-directional margin, each modal feature of the same pedestrian target is compressed compactly, making decision boundaries clearer. We then relax the constraints to:
$0.5D^{2}(A_{img},A_{txt})+\mathcal{M}_{1}<0.5D^{2}(A_{img},P_{txt}),$ (3)
$0.5D^{2}(A_{img},P_{txt})+\mathcal{M}_{2}<0.5D^{2}(A_{img},N_{txt}).$ (4)
Subsequently, we change the above pairwise constraints into soft forms for better convergence following [33].
$\mathcal{L}^{Pull}_{match}=log\big{[}1+\sum\limits^{K}_{k=1}exp(\alpha(\bar{v}_{i}^{c}\bar{t}_{k}^{c}-\bar{v}_{i}^{c}\bar{t}_{i}^{c}+\mathcal{M}_{1}))\big{]},$ (5)
$\mathcal{L}^{Push}_{match}=log\big{[}1+\sum\limits^{K}_{k=1}\sum\limits^{J}_{j=1}exp(\alpha(\bar{v}_{i}^{c}\bar{t}_{j}^{c}-\bar{v}_{i}^{c}\bar{t}_{k}^{c}+\mathcal{M}_{2}))\big{]},$ (6)
where $\bar{v}^{c}$ and $\bar{t}^{c}$ are CLS token features after normalizing, $K$ and $J$ denote the number of positive and negative samples in this batch, respectively. $\alpha$ is a scale parameter. As a result, the image-to-text matching part of sew calibration loss is formulated as below:
$\mathcal{L}^{I2T}_{match}=\mathcal{L}^{Pull}_{match}+\mathcal{L}^{Push}_{match}.$ (7)
The constraints in Eq. 2 can also be used to impose a classification loss in a similar way. As there is no difference for perfect positive samples in the classification task, we omit to constrain Eq. 1. Formally, the loss for our person ID classification part is as follows [16]:
$\mathcal{L}^{I2T}_{id}&=\frac{1}{n}\sum\limits_{i=1}^{n}-log\big{(}\frac{e^{(\alpha(s_{y_{i},i}-\mathcal{M}_{2}))}}{e^{(\alpha(s_{y_{i},i}-\mathcal{M}_{2}))}+\sum\limits_{j\neq{y_{i}}}e^{(\alpha(s_{y_{j},i}))}}),\end{aligned}$ (8)
$s_{i,j}=\omega^{T}_{i}\hat{v}_{j},\quad\hat{v}_{i}=(v^{c}_{i})^{T}\bar{t}^{c}_{i}\cdot\bar{t}^{c}_{i},$ (9)
where $n$ is batch size, and $\omega$ represents the classification weight after normalization. $\hat{v}_{i}$ can be explained as the projection of image representation $v^{cls}_{i}$ onto the normalized text representation $\bar{t}^{c}_{i}$.
$\mathcal{L}^{T2I}_{match}$ and $\mathcal{L}^{T2I}_{id}$ are in the same form as above, but the change is focused on the text-to-image. Both matching and classification loss have identical decision boundaries. Equipped with our proposed cross-modal constraints, the sew calibration loss can effectively reduce the gap between image and text feature distributions. Although the margin restrictions allow our model to learn better cross-modal representations, using a fixed margin $\mathcal{M}$ in all cases may not be flexible enough. A large margin constraint makes model learning difficult, while too small a margin does not impose a significant constraint. An adaptive margin guided by quality can be more effective. In TBPS, the texts come from the annotator’s descriptions of the images. Images are expected to have complete information, while texts have varying amounts of information. Thus, we adjust the margin value based on the quality of the text description. We argue that a less informative caption (i.e., a shorter caption) needs a smaller margin as a looser constraint. Based on this, we compute the adaptive margin for each image-text pair according to its text total tokens length $\mathcal{T}_{i}$:
$\mathcal{M}_{i}=\mathcal{M}_{min}+\frac{(\mathcal{M}_{max}-\mathcal{M}_{min})\cdot(\mathcal{T}_{i}-\mathcal{T}_{min})}{\mathcal{T}_{max}-\mathcal{T}_{min}},$ (10)
where $\mathcal{M}_{max}$ and $\mathcal{M}_{min}$ are upper and lower bounds of margins. $\mathcal{T}_{max}$ and $\mathcal{T}_{min}$ are bounds of the captions length, respectively. We set $\mathcal{T}_{max}$ and $\mathcal{T}_{min}$ according to the different dataset captions length distributions. After that, we utilize $\mathcal{M}_{i}$ to replace the fixed margin $\mathcal{M}$ above. We simply set $\mathcal{M}_{1}=\mathcal{M}_{2}=\mathcal{M}_{i}$.
3.2 Masked Caption Modeling Loss
TBPS is a fine-grained cross-modal task, which means only caption-level discrimination is not enough. If the textual captions of two persons differ in a few words, a TBPS method can not retrieve a specific person without word-level discrimination. Although there are many works to establish word-level discrimination capacities, such methods are complex and limited, hindering performance boost. To solve this issue, we propose masked caption modeling to establish detailed image-text relationships. Furthermore, by utilizing MCM, our framework can perform more generic cross-modal learning.
Inspired by [6, 35], we add a masked prediction task on the text branch. Concretely, this loss is based on a cross-modal decoder architecture. We mask a portion of text tokens and replace these masked tokens with a learnable token vector. The text encoder inputs these text tokens and outputs the corresponding text features. The cross-modal decoder $f_{cd}$ learns to maximize the conditional likelihood of the masked text feature $t_{n}$ under latent image feature sequence $\{v_{i}\}$ and text feature sequence $\{t_{i}\}$:
$\mathcal{L}_{mcm}=-\sum_{n=1}^{N}\text{log }P_{\theta}(t_{n}|\{t_{i}\},\{v_{i}\}),$ (11)
where $N$ is the total masked token numbers in a caption.
Input:
Image $I$ and text $T$; A batch of $n$ paired $\mathbb{G}=\{(I^{1},T^{1}),(I^{2},T^{2}),\cdots,(I^{n},T^{n})\}$
Output: Training loss $(\mathcal{L}_{sew},\mathcal{L}_{mcm})$  or  $(I_{cls},T_{cls})$
foreach $T^{i}$ in set $\mathbb{G}$ do
      $T^{i}\leftarrow\text{mask}(T^{i})$;
end foreach
for $i\leftarrow 1$ to $\mathbb{G}$ do
       $(I^{i}_{cls},I^{i}_{1},\cdots,I^{i}_{c})\leftarrow ViT(I^{i})$;
       $(T^{i}_{cls},T^{i}_{1},\cdots,T^{i}_{c})\leftarrow Bert(T^{i})$;
       if Training Stage then
             $(I^{i}_{cls},T^{i}_{cls})\leftarrow BatchNorm(I^{i}_{cls},T^{i}_{cls})$;
             $I^{i}_{context},T^{i}_{context}\leftarrow(I^{i}_{1},\cdots,I^{i}_{c}),(T^{i}_{1},\cdots,T^{i}_{c})$;
             $T^{i}_{context}\leftarrow Attn(T^{i}_{context})$;
             $T^{i}_{cross}\leftarrow CrossAttn(T^{i}_{context},I^{i}_{context})$;
             $\mathcal{L}_{sew}^{i}\leftarrow SewCalibration(I^{i}_{cls},T^{i}_{cls})$;
             $\mathcal{L}_{mcm}^{i}\leftarrow MCM(T^{i}_{cross},T^{i})$;
      else
             $\{(I^{1}_{cls},T^{1}_{cls}),(I^{2}_{cls},T^{2}_{cls}),\cdots,(I^{i}_{cls},T^{i}_{cls})\}\leftarrow(I_{cls},T_{cls})$;
       end if
end for
if Training Stage then
       return $(\mathcal{L}_{sew},\mathcal{L}_{mcm})$;
else
       return $(I_{cls},T_{cls})$;
end if
Algorithm 1 Training Procedure of SCMM
As shown in Fig. 2, the cross-modal decoder part, the decoder $f_{cd}$ takes both unmasked text tokens and masked tokens in their original order as the input. The multi-head self-attention [18] first encodes the text features as $Q_{t},K_{t},V_{t}$, while the cross self-attention further improves the text features by taking into account the encoded image features as $K_{i},V_{i}$ for visual context. The final linear projection layer has the same number of output channels as the text vocabulary and computes the cross entropy loss between the reconstructed and original words only on masked text tokens. It should be noted that $f_{cd}$ is only used during training and not during inference.
3.3 Total Loss
The total loss function of SCMM integrates our two complementary fusion mechanisms to achieve comprehensive cross-modal alignment and correspondence learning. Our unified optimization objective combines the sew calibration loss and the masked caption modeling loss to jointly address both global alignment and fine-grained correspondence challenges in cross-modal information fusion. The sew calibration loss $\mathcal{L}_{sew}$ encompasses bidirectional matching and classification components that ensure symmetric cross-modal alignment:
$\mathcal{L}_{sew}=\mathcal{L}_{match}^{I2T}+\mathcal{L}_{match}^{T2I}+\mathcal{L}_{id}^{I2T}+\mathcal{L}_{id}^{T2I},$ (12)
where the image-to-text and text-to-image matching losses establish cross-modal correspondence constraints, while the identification losses enforce discriminative learning for person-specific features. Importantly, the bidirectional formulation ensures that both modalities are effectively aligned in the shared embedding space through symmetric constraints. Subsequently, the complete loss function combines both components through weighted summation:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{sew}+\lambda_{2}\mathcal{L}_{mcm},$ (13)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters balancing global alignment and fine-grained correspondence learning. Through this joint optimization, the framework simultaneously learns compact cross-modal representations while capturing detailed visual-textual relationships for discriminating similar individuals. The complementary loss components create synergistic effects ensuring comprehensive information fusion, addressing both inter-modal gap reduction and fine-grained correspondence modeling. Regarding computational complexity, SCMM maintains efficient processing through its streamlined dual-encoder architecture. Training complexity is $O(n\cdot d^{2})$ where $n$ is batch size and $d$ is feature dimension, dominated by cross-attention operations in the MCM decoder. During inference, only CLS token embeddings are utilized, achieving $O(d)$ complexity per query-gallery comparison. In contrast, the method contrasts favorably with attention-explicit methods like ACSA [26] requiring $O(n^{2}\cdot d)$ complexity due to full cross-modal attention during inference. Our training-only decoder design ensures high-speed retrieval while maintaining sophisticated cross-modal learning capabilities.
4 Experiments
4.1 Datasets and Evaluation Metric
We evaluate SCMM on three benchmarks: CUHK-PEDES [8], ICFG-PEDES [20], and RSTPReid [13]. CUHK-PEDES is the first large-scale benchmark for text-based person search tasks. As the foundational dataset, it contains 40,206 images of 13,003 person IDs collected from five person re-identification datasets. Each image has two different textual captions with an extensive vocabulary, and the average sentence length is 23.5. The testing set comprises 3,074 images and 6,148 descriptions of 1,000 persons. ICFG-PEDES contains 54,522 images of 4,102 persons. For each image, the corresponding description has an average length of 37 words. The testing set consists of 19,848 image-text pairs of 1,000 persons. RSTPReID contains 20,505 images of 4,101 persons, with each pedestrian having five images. It is divided into a training set with 3,701 persons, a validation set with 200 persons, and a testing set with 200 persons. For our evaluation metric, we report the Rank $k$ ($k$=1, 5, 10) text-to-image accuracy, which is commonly used in previous works to evaluate text-based person search. Given a textual description as the query, if the top-$k$ retrieved images contain any person corresponding to the query, we consider it a successful person search.
4.2 Implementation Details
In our visual-textual dual-encoder, we extract visual representations using the ViT-Base pre-trained on ImageNet [38]. The images are resized to 224 $\times$ 224 pixels. For textual representations, we use the BERT-Base-Uncased model pre-trained on the Toronto Book Corpus and Wikipedia. The representation dimension is set to 768, and the feature sequence lengths are set to 197 and 100, respectively. During the training phase, we use a batch size of 64 and train for 60 epochs. We use Adam optimizer with an initial learning rate of 0.001. To augment our data, we apply a random horizontal flipping operation, and we use a mask ratio of 0.1 for randomly masking text tokens. The minimum and maximum textual information length boundaries $\mathcal{T}_{min}$ and $\mathcal{T}_{max}$ are set to 20-60, 25-65, and 22-60 according to the caption length distributions of CUHK-PEDES, ICFG-PEDES, and RSTPReID, respectively. The bounds of the upper and lower margin $\mathcal{M}$ are set to 0.4 and 0.6, and the scale parameter $\alpha$ is set to 32. For each loss in the total loss function, the balance factors $\lambda_{1}$ and $\lambda_{2}$ are set equal to 1. During the testing phase, we apply a re-ranking post-processing approach to improve search performance following NAFS [14]. We conduct the experiments on four NVIDIA 2080Ti GPUs using PyTorch.
4.3 Comparison with State-of-the-art Methods
Table 1: Comparison with SOTA methods on the CUHK-PEDES dataset. Rank1, Rank5, and Rank10 accuracies (%) are reported. Bold number represents the best score. R and C denote the re-ranking post-processing operations and image encoder pre-trained on CLIP model.
Method Rank1 Rank5 Rank10
GNA-RNN [8] 19.05 - 53.64
CMPM/C [16] 49.37 - 79.27
Dual-path [15] 44.40 66.26 75.07
ViTAA [19] 55.97 75.84 83.52
VP Net [27] 58.83 81.25 86.72
SUM [39] 59.22 80.35 87.60
CLIP [28] 60.67 81.99 88.87
DSSL [13] 59.98 80.41 87.56
MGEL [23] 60.27 80.01 86.74
SSAN [20] 61.37 80.15 86.73
TestReID [40] 64.08 81.73 88.19
ACSA [26] 63.56 81.40 87.70
IVT [37] 64.00 82.72 88.95
SAFA Net [22] 64.13 82.62 88.40
TIPCB [21] 64.26 83.19 89.10
CAIBC [12] 64.43 82.87 88.37
AXM Net [41] 64.44 80.52 86.77
CFine [25] 65.07 83.01 89.00
ISANet [5] 63.92 82.15 87.69
SCMM 67.71 84.57 89.44
SCMM+R 71.09 86.78 91.23
SCMM+C 69.61 86.01 90.90
SCMM+R+C 73.81 88.89 92.77
4.3.1 Results on CUHK-PEDES.
Table 1 compares our framework and previous methods on CUHK-PEDES. It can be observed that our proposed SCMM can outperform all previous methods by a large margin. Compared to the work CFine [25], SCMM achieves 67.71% (+2.64%), 84.57% (+1.56%) and 89.44% (+0.44%) on Rank1, Rank5 and Rank10 without re-ranking. The state-of-the-art results on the CUHK-PEDES show the effectiveness of SCMM. With the help of re-ranking, SCMM shows an incremental boost to get a 71.09% Rank1 score. Furthermore, with a better image encoder pre-training on CLIP, SCMM achieves 73.81%, 88.89% , and 92.77% on three metrics, respectively. Overall, the consistent improvements show the scalability of SCMM across better pre-trained models and extra post-processing operations.
Table 2: Quantitative results on the ICFG-PEDES dataset.
Method Rank1 Rank5 Rank10
CMPM/C [16] 43.51 65.44 74.26
Dual-path [15] 38.99 59.44 68.41
ViTAA [19] 50.98 68.79 75.78
SSAN [20] 54.23 72.63 79.53
CLIP [28] 53.96 73.69 80.43
TIPCB [21] 54.96 74.72 81.89
IVT [24] 56.04 73.60 80.22
CFine [25] 55.69 72.72 79.46
CFine+C [25] 60.83 76.55 82.42
IRRA [34] 63.46 80.25 85.82
TP-TPS [11] 60.64 75.97 81.76
ISANet [5] 57.73 75.42 81.72
CANC [2] 60.52 78.36 84.13
SCMM 60.20 75.97 81.78
SCMM+R 62.17 85.74 89.67
SCMM+C 62.29 77.15 82.52
SCMM+R+C 64.25 86.95 90.70
Table 3: Quantitative results on the RSTPReid dataset.
Method Rank1 Rank5 Rank10
CLIP [28] 50.10 76.10 84.95
DSSL [13] 32.43 55.08 63.19
SUM [39] 41.38 67.48 76.48
SSAN [20] 43.50 67.80 77.15
IVT [24] 46.70 70.00 78.80
ACSA [26] 48.40 71.85 81.45
CFine [25] 45.85 70.30 78.40
CFine+C [25] 50.55 72.50 81.60
TP-TPS [11] 50.65 72.45 81.20
SCMM 50.75 74.20 81.70
SCMM+R 55.35 77.30 84.25
SCMM+C 51.95 73.50 82.45
SCMM+R+C 57.35 77.50 85.50
4.3.2 Results on ICFG-PEDES and RSTPReid.
We also utilized other benchmarks to validate SCMM’s performance and generalization. The ICFG-PEDES and RSTPReid datasets are more challenging compared to CUHK-PEDES, and our method significantly outperformed all state-of-the-art methods on these two datasets by a large margin, as reported in Tables 2 and 3. Compared with the state-of-the-art results [25] on ICFG-PEDES, SCMM achieved significant improvements, with scores of 60.20% (+4.51%), 75.97% (+3.25%), and 81.78% (+2.32%) on Rank1, Rank5, and Rank10, respectively. On the RSTPReid dataset, we achieved scores of 50.75% (+4.90%), 74.20% (+3.90%), and 81.70% (+3.30%) on the three metrics. With re-ranking post-processing, we were able to achieve scores of 62.17% and 55.35% on Rank1 for the two benchmarks, respectively.
We note that re-ranking also brings a significant boost, as our model learns clear and compact cross-modal key information patterns, and re-ranking can retrieve the correct feature neighbors more effectively. Moreover, by utilizing the CLIP pre-trained image model as an encoder, we achieved even better results, with scores of 64.25%, 86.95%, and 90.70% on ICFG-PEDES, and scores of 57.35%, 77.50%, and 85.50% on RSTPReid. Consequently, the two challenging benchmark results demonstrate the robustness and scalability of SCMM.
4.4 Ablation Studies
Table 4: Performance comparisons of different components. Res and ViT denote ResNet and Vision Transformer as image encoders, respectively. Sew-F and Sew-A refer to the fixed and adaptive margin versions of the sew calibration loss, respectively. Model 2 is our baseline.
ID Res ViT Sew-F Sew-A MCM Rank1 Rank5 Rank10
1 60.39 80.22 86.53
2 62.29 81.25 87.22
3 65.41 82.56 88.01
4 66.02 83.33 88.61
5 65.16 81.95 87.93
6 67.71 84.57 89.44
4.4.1 Analysis of Model Components
To fully validate the performance of the different components in SCMM, we demonstrate the contributions of each part on CUHK-PEDES, as shown in Table 4. Model 1 and Model 2 show the results using ResNet [42] and ViT [37] as the image encoder, respectively, without the Sew Calibration and MCM losses. We utilize Model 2 as our baseline, and CMPM and CMPC [16] are used as loss functions in the baseline experiment. First, we compare ResNet-50 and ViT-Base as the image encoder and observe a performance improvement of 1.90%, 1.03%, and 0.69% on Rank1, Rank5, and Rank10, respectively. Based on this, we adopt ViT as the image encoder in the following experiments. We then validate the sew calibration and MCM losses compared to the baseline. From Models 2-4, we observe that sew calibration loss brings marginal improvement, and a common distribution can provide a better basis for optimizing the embedding space for fine-grained cross-modal recognition. Moreover, Model 3 represents the sew calibration loss with the fixed margin 0.5, while Model 4 represents the sew calibration loss with adaptive margins. We observe that Model 4 achieves better results than a fixed manual margin, no matter how we tune the margin value, which demonstrates the effectiveness of adaptive constraints. We also find that MCM gives another substantial improvement to the baseline from Model 2 to Model 5, indicating that the text-image detail mining capability is critical. Notably, when the two components are used jointly in Model 6, SCMM continues to improve and outperform the baseline by 4.57%, 2.99%, and 2.37% on the three metrics, respectively. Importantly, the results demonstrate that obtaining well-distributed image-text features in the common embedding space is essential for reducing the cross-modal gap. The consistent improvement in each component of the method demonstrates our effectiveness.
Figure 4: Comparison of singular values for image-text embedding features across CUHK-PEDES (a-b) and ICFG-PEDES (c-d) datasets. (a) and (c) depict the distribution of singular values, where a smaller inter-line gap indicates a closer cross-modal distribution. (b) and (d) present the logarithmic difference in singular values between the baseline and our approach for CUHK-PEDES and ICFG-PEDES, respectively.
4.4.2 Impact of Sew Calibration Loss for Reducing Cross-modal Gap
Benefiting from the sew calibration loss, which reduces the cross-modal gap, the learned representation distributions are closer than with the basic loss. To demonstrate this, we illustrate a comparison of singular value decomposition on cross-modal representations, inspired by [3]. Specifically, Fig. 4(a) and (c) present the baseline and SCMM’s singular value decomposition for the text and image modalities. We computed the distance between these two modalities at specific singular values and reflected them in Fig. 4(b) and (d). Intuitively, the smaller the distance between different modality features, the closer the distribution learned by the model. We find that the sew calibration loss performs better representation distributions in the common embedding space. It ensures closer cross-modal representation distributions and a smaller inter-modal gap than the baseline.
4.4.3 Impact of Masked Caption Modeling
The masked caption modeling operation in fine-grained cross-modal interaction is critical in our framework. It comprises tokens masking, attention module, and masked tokens prediction. As shown in Table 5, we explore the effectiveness of MCM components on the CUHK-PEDES. First, only masking on the input text tokens without a reconstruction task behaves like a random erase text augmentation. Remarkably, the augmentation can already bring +1.36% marginal improvement and reach 63.65% on Rank1. It shows that the details provided by word tokens are helpful for fine-grained recognition.
Next, if we only utilize attention mechanism to enhance cross-modal representations without the mask caption modeling, it can achieve 63.83% on Rank1. We explain this improvement as the cross-attention brings details from sequence tokens to the CLS token. Meanwhile, image features also contribute a lot to this process.
On the other hand, we also design an experiment of mask caption modeling without the cross attention. The decoder $f_{cd}$ directly predicts all masked tokens from dual-encoder outputs. In this experiment, we observe 64.28% on Rank1. It proves the reconstruction tasks in the decoder help guide the text encoder to learn richer and more refined representations. We also notice that with all three components, the MCM achieves 65.16% on Rank1. With the help of cross attention from image features, the CLS token ensembles all those rich information.
Table 5: MCM components on the CUHK-PEDES dataset.
Mask Attention Caption Modeling Rank1 Rank5 Rank10
63.65 81.84 87.51
63.83 80.87 86.35
64.28 81.97 87.75
65.16 81.95 87.93
4.4.4 Analysis of Generalisability Validation
Table 6: Performance analysis for cross-domain validation.
      CUHK $\Rightarrow$ ICFG       Rank1       Rank5       Rank10
      Baseline       30.79       49.10       57.47
      Baseline+MCM       31.13       49.52       58.07
      ICFG $\Rightarrow$ CUHK       Rank1       Rank5       Rank10
      Baseline       22.19       40.55       50.52
      Baseline+MCM       23.78       42.41       52.11
      CUHK $\Rightarrow$ RSTP       Rank1       Rank5       Rank10
      Baseline       37.40       63.40       74.35
      Baseline+MCM       39.25       63.95       74.55
      RSTP $\Rightarrow$ CUHK       Rank1       Rank5       Rank10
      Baseline       10.02       22.95       31.04
      Baseline+MCM       10.69       25.20       33.93
To better understand the contributions of MCM in our framework, we conduct a domain generalization analysis on the three benchmarks to demonstrate our generalization performance. In Table 6, CUHK $\Rightarrow$ ICFG and CUHK $\Rightarrow$ RSTP indicate using the model trained on the CUHK-PEDES dataset to infer their test sets and vice versa. Compared to the baseline, we can observe a performance improvement in the three metrics for ICFG-PEDES and RSTPReid. The fine-grained features mined by MCM are more resistant to overfitting. The improvement in domain generalization shows the capability of our framework in learning generic cross-modal information, which is essential to solve the fine-grained modal heterogeneity.
4.4.5 Effect of Margin Parameters
In Sec.  3.1, we introduce the adaptive margins to constrain cross-modal learning using Sew Calibration loss with constraints. We find that compared to the manually fixed parameters (SEW-f), our adaptive margins (SEW-a) produce better results as shown in Fig. 5, where we investigate the effect of manual margin in different settings on the CUHK-PEDES [8] on Rank1 accuracy. The best performance of 65.41% is achieved when $\mathcal{M}=0.5$, while the worst model performance of 62.29% is observed when $\mathcal{M}=0$. We note that the performance of the model first improves as the margin increases, which is due to our bi-directional margin compressing the same pedestrian target compactly. However, we find that the model performance does not continue to improve as the margin increases, and we observe an inflection point around margin 0.5. When the margin is too large, the tight constraints make the model too hard to train.
Therefore, it is critical to utilize quality-guided adaptive margins according to the textual information. In this regard, we set the upper and lower margin bounds ($\mathcal{M}_{min}=0.4$ and $\mathcal{M}_{max}=0.6$). In this range, our Sew Calibration loss with adaptive constraints obtains the best performance, which a manually fixed margin cannot achieve, regardless of how the margin value is tuned.
4.4.6 Effect of Mask Ratio
Inspired by Masked Language Modeling [36], we propose the MCM loss which leverages a masked captions prediction task to establish detailed and generic relationships between textual and visual parts. Our method requires setting the mask ratio in the MCM. To investigate the effect of the mask ratio on model performance and generalization learning ability, we conducted ablation experiments, and the results are shown in Table 7. When the mask ratio is set to 0, which is equivalent to the baseline, the performance is the worst. Compared to the baseline, our method with a 0.1 mask ratio achieves the best results, with scores of 65.16%, 81.95%, and 87.93% on Rank1, Rank5, and Rank10, respectively.
Table 7: Ablation results of mask ratio on the CUHK-PEDES.
Baseline+MCM Rank1 Rank5 Rank10
Mask = 0 62.29 81.25 87.22
Mask = 0.1 65.16 81.95 87.93
Mask = 0.3 64.97 81.38 87.71
Mask = 0.5 63.78 81.11 87.43
The experiments are conducted on the CUHK-PEDES dataset using the Baseline+MCM method. As shown in these results, we find that the model achieves the best performance when the mask is set to 0.1, and then the model performance gradually decreases as the mask ratio increases. The main reason is that the information provided in the annotated caption text is very accurate and free of redundant information. Therefore, as the mask ratio expands, more and more key information may be obscured, causing the model to fail to learn enough semantic information to complete the construction of the image-text relationship.
Figure 5: Effect of the manual fixed margin parameters setting of our Sew Calibration loss in terms of Rank1 accuracy on the CUHK-PEDES.
4.5 Qualitative Analysis
4.5.1 Visualization of Attention Map
Figure 6: Visualization of attention maps from baseline and SCMM on the CUHK-PEDES. We present the total caption-level results and the fine-grained word-level results, respectively. (1) Baseline caption-level results, (2) Our caption-level results, (3) & (4) Our word-level results. Best viewed in color.
We visualize attention maps on the CUHK-PEDES test dataset to demonstrate model capability in learning image-text correspondence. As shown in Fig. 6, compared to the baseline, we can observe that the visualization results obtained by SCMM are more apparent and refined. We conduct the word-level visualization to validate further the ability to perform fine-grained interaction. We select several keywords in the caption description, e.g., bag, shoes as items, colors, and clothing as modifiers. For example, the visualization results of IDs 8491 and 10648 do not focus on useful detailed information. The key messages in the two images are a man riding bike with a backpack and a woman wearing a sleeveless white dress. We can observe that SCMM successfully captures the key detail information compared to the baseline. From the word-level results, we can also observe that SCMM learns the critical parts in the cross-modal correspondence.
Figure 7: Presentation of cross-modal representation distributions on CUHK-PEDES test dataset. Different colors correspond to the different target ID. Best viewed in color.
Figure 8: Top 5 retrieval results on CUHK-PEDES. Matched target person images are marked with red rectangles. and subcaptions show the text queries used for retrieval.
4.5.2 Visualization of Image-text Feature Distribution
SCMM is capable of learning better image-text representation distributions with fine-grained image-text interaction and adaptive constraints to reduce the cross-modal gap. To demonstrate this, we randomly selected some pedestrians and extracted their image-text global features, then mapped them to two dimensions for visualization in Fig. 7(a) and (b). For instance, we can take IDs 5532 and 11800 as examples. We can observe that the image-text features of the same ID distribution are compressed more compactly, and the boundaries between them are more apparent than the baseline results, which demonstrates the effectiveness of SCMM in mitigating the cross-modal gap.
4.5.3 Visualization of Top 5 Search Qualitative Results
To evaluate the performance of our method, we randomly selected several sets of search results from our method and visualized them in Fig. 8. For each target image, we show the top 5 most similar search results, with the similarity scores after re-ranking processing displayed above each image. Correct search results are indicated by red rectangles, and the number in the last column shows how many pedestrian targets with the same ID have been recalled. Taking ID 5219 for an example, our method correctly retrieved all three target images, with similarity scores of 1.2563, 1.2100, and 0.9775, respectively. For the last two unmatched images, our method gave similarity scores of 0.5174 and 0.5011, respectively. Although these images have some similarities with the positive samples, they do not match the text description in other details, and our method learned to differentiate them. The last row in Fig. 8 shows a hard case, where the man in the query image has extremely similar attributes to several other pedestrians. Despite, our method is able to correctly rank 2/3 of the right person images as the top 2 search results.
5 Conclusion
In this work, we introduce SCMM, a novel AI framework for text-based person search that addresses fundamental challenges in cross-modal learning. Our method introduces two innovative components: sew calibration loss with adaptive margins and masked caption modeling loss that establishes fine-grained correspondences through transformer-based architectures. These complementary strategies bridge the inter-modal gap while capturing detailed relationships essential for person identification. Extensive experimental validation demonstrates SCMM’s superior performance, achieving state-of-the-art results of 73.81%, 64.25%, and 57.35% Rank1 accuracy on CUHK-PEDES, ICFG-PEDES, and RSTPReID respectively. SCMM’s streamlined dual-encoder architecture prioritizes computational efficiency, making it suitable for real-time applications. Future work will explore incorporating additional modalities and developing more sophisticated adaptive mechanisms for multimodal learning in person re-identification tasks.
CRediT authorship contribution statement
Jing Liu: Conceptualization, Methodology, Formal analysis, Writing – original draft, Writing – review & editing, Visualization. Donglai Wei: Methodology, Software, Validation, Formal analysis, Investigation, Writing – review, Project administration & editing. Yang Liu: Conceptualization, Methodology, Investigation, Writing – review & editing, Visualization. Sipeng Zhang: Writing – review & editing, Funding acquisition. Tong Yang: Writing – review & editing, Funding acquisition. Wei Zhou: Conceptualization, Writing – review & editing. Weiping Ding: Methodology, Writing – review & editing. Victor C. M. Leung: Supervision, Writing – review & editing, Project administration.
Data availability
Data will be made available on request.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used generative AI and AI-assisted technologies for writing enhancement and language polishing to improve the clarity and readability of the manuscript. After using these tools, the authors have reviewed and edited the content as needed and take full responsibility for the content of the publication.
Acknowledgements
This work was supported in part by the Mitacs Elevate under Grant No. IT44479; the Guangdong Pearl River Talents Recruitment Program under Grant No. 2019ZT08X603; and the Guangdong Pearl Rivers Talent Plan under Grant No. 2019JC01X235. We sincerely thank all the editors and anonymous reviewers for their careful work and thoughtful suggestions that have helped improve this paper substantially.